A. List the members of the following sets

1. {x| x is real numbers and x2 = 1}

2. {x| x is an integer and -4 < x ≤ 3}


B. Use set builder notation to give description of each of these sets.

1. {a, e,i ,o, u}

2. {=2, -1, 0, 1, 2}

C. Let A= (a, b, c), B = (x, y) and C = (0, 1)

Find:

1. A U C

2. C x B

3. B – A

4. (A ∩ C) U B


D. Find these terms of the sequence (An}, where An = 2(3)n + 5

1. A0

2. A5

3. A3

4. 8th term

5. 2nd term

6. Sum of the sequence


E. Given the following set:

2. X = {-1, 0, 1, 2, 3, 4, 5} defined by the rule (x, y) ∈R if x ≤ y

F. List the elements of R

G. Find the domain of R

H. Find the range of R

I. Draw the digraph

J. Properties of the Relation


1
Expert's answer
2021-03-18T15:47:47-0400

A. (1). {"-1,1" }

(2). {"-3,-2,-1,0,1,2,3" }


B. (1). {"x|x" is a vowel}

(2). {"x|x" is an integer and "-2\leq x\le 2" }


C.(1) "A\cup C =" {"a,b,c,0,1" }

(2) "C\u00d7B=" {"(0,x),(0,y),(1,x),(1,y)" }

(3) "B-A=" { "x,y" }

(4) "(A\cap C)\cup B=" "\phi \cup B =B=" {"x,y" } [ Since "A" and C have no common element therefore "A\cup C=\phi" ]


D. Given sequence is "A_n=2^3.n+5"

(1) "A_0=5"

(2) "A_5=( 2^3.5+5)=45"

(3) "A_3=( 2^3.3+5)=29"

(4) 8 th term "=A_7=( 2^3.7+5)=61"

(5) 2nd term "=A_1=( 2^3.1+5)=13"

(6) Sum of the sequence "=S_n=(A_1+A_2+A_3+......+A_n)"

"=2^3(1+2+3+....+n)+5n"

"=2^3.\frac{n.(n+1)}{2}+5n"

"=n[4n+9]"


E. (F). "R=" {"(-1,-1),(-1,0),(-1,1),(-1,2),(-1,3),(-1,4),(-1,5),(0,0),(0,1),(0,2),(0,3),(0,4),(0,5),(1,1)"

"(1,2),(1,3),(1,4),(1,5),(2,2)(2,3),(2,4),(2,5),(3,3),(3,4),(3,5),(4,4),(4,5),(5,5)" }


(G)Domain of "R" "=" {"-1,0,1,2,3,4,5" }

(H) Range of "R=" {"-1,0,1,2,3,4,5" }

(I) Digraph of R given below



(J) As we seen above that their is a loop in each point. Therefore "R" is reflexive.

But not symmetric as "(0,2)\in R" but "(2,0)\notin R" ,

Also the relation is transitive .